# Resp-Agent: An Agent-Based System for Multimodal Respiratory Sound Generation and Disease Diagnosis

### v1

@@ Abstract
Deep learning-based respiratory auscultation is currently hindered by two fundamental challenges: (i) inherent information loss, as converting signals into spectrograms discards transient acoustic events and clinical context; (ii) limited data availability, exacerbated by severe class imbalance. To bridge these gaps, we present Resp-Agent, an autonomous multimodal system orchestrated by a novel Active Adversarial Curriculum Agent (Thinker-A2CA). Unlike static pipelines, Thinker-A2CA serves as a central controller that actively identifies diagnostic weaknesses and schedules targeted synthesis in a closed loop. To address the representation gap, we introduce a Modality-Weaving Diagnoser that weaves EHR data with audio tokens via Strategic Global Attention and sparse audio anchors, capturing both long-range clinical context and millisecond-level transients. To address the data gap, we design a Flow Matching Generator that adapts a text-only Large Language Model (LLM) via modality injection, decoupling pathological content from acoustic style to synthesize hard-to-diagnose samples. As a foundation for these efforts, we introduce Resp-229k, a benchmark corpus of 229k recordings paired with LLM-distilled clinical narratives. Extensive experiments demonstrate that Resp-Agent consistently outperforms prior approaches across diverse evaluation settings, improving diagnostic robustness under data scarcity and long-tailed class imbalance. Our code and data are available at https://github.com/zpforlove/Resp-Agent.
@@ 1 Introduction
The first challenge is a unimodal representational bottleneck. Audio models often convert signals into mel-spectrograms for image-style CNNs (Bae et al., 2023; He et al., 2024), which discards phase and blurs fine temporal structure, obscuring transient events such as crackles(Paliwal et al., 2011). Conversely, text-only models capture electronic health record (EHR) context but lack objective acoustic evidence, limiting discrimination between conditions with similar narratives but distinct auscultatory patterns. Without deep multimodal fusion, performance and reliability saturate.
@@ 2 Related Work
Respiratory Sound Classification (RSC) has largely relied on audio-only pipelines trained on small, single-source datasets, which limits out-of-domain generalization. Standard approaches utilize pretrained backbones like PANNs (Kong et al., 2020) and AST (Gong et al., 2021) to mitigate data scarcity. The OPERA benchmark has begun to close the data gap via domain-specific pretraining (Zhang et al., 2024a); however, most systems remain constrained by single-modality supervision and in-distribution evaluation. Our work departs from this paradigm in two key ways. (i) Multimodal fusion. We weave EHR-style textual tokens and acoustic tokens within a long-context Transformer. Unlike RespLLM (Zhang et al., 2024b), which feeds concatenated modality tokens through dense full attention, we introduce Strategic Global Attention with sparse audio anchors, drawing on ideas from efficient Transformers (Beltagy et al., 2020; Zaheer et al., 2020) to route clinical context to transient acoustic events at sub-quadratic cost. We evaluate on source-disjoint splits to stress-test generalization under realistic distribution shifts (Koh et al., 2021) and label imbalance (Johnson and Khoshgoftaar, 2019). (ii) Targeted augmentation. Recent generative models enable high-fidelity audio synthesis (Borsos et al., 2023; Lipman et al., 2022; Liu et al., 2022; Peebles and Xie, 2023), but existing augmentation strategies are typically untargeted, relying on generic perturbations such as SpecAugment (Park et al., 2019) or unconditional generation (Kim et al., 2023). We introduce Resp-Agent, a closed-loop system in which an LLM-based Thinker-A2CA diagnoses model failures and requests condition-controlled synthesis from a flow-matched generator, turning augmentation into a precise instrument for adversarial edge-case creation and distribution balancing.
@@ 3 Resp-229k: A Large-Scale, Multi-Source, Cross-Domain Benchmark
We introduce Resp-229k to address the scarcity of multimodal supervision and the lack of robust cross-domain evaluation in respiratory sound analysis. Unlike existing datasets, Resp-229k provides paired audio with standardized clinical summaries, converting diverse metadata into a format suitable for multimodal modeling. We also establish a strict out-of-domain evaluation protocol to explicitly test model generalization. The dataset comprises 229,101 quality-controlled samples sourced from five public databases, categorized into 16 classes (15 conditions and 1 control).
@@ 4 Resp-Agent: An LLM-Orchestrated Loop for Unified Diagnosis and Controllable Synthesis
The overall architecture of Resp-Agent is depicted in Figure 1. Given the paired text–audio supervision and the cross-domain split established by Resp-229k, Resp-Agent is designed as a centrally planned, compute-aware multi-agent system that integrates standalone audio and NLP modules into a closed loop. A compute-efficient planner, Thinker-A2CA (DeepSeek-V3.2-Exp; (Guo et al., 2025)), performs semantic intent parsing and plan–execute tool routing using structured arguments (audio paths, EHR tables, and target classes), enforcing deterministic I/O schemas and emitting concise, instrumented logs. Beyond dispatch, the controller reuses model rationales, error profiles, and calibrated confidence to bias subsequent actions (e.g., targeted synthesis for failure modes), thereby coupling data generation and diagnosis under tight accelerator budgets without compromising coverage or reproducibility. The Thinker coordinates two task-specific agents detailed below: a Generator (Section 4.1) that synthesizes controllable respiratory audio via modality injection and conditional flow matching, and a Diagnoser (Section 4.2) that fuses clinical narratives with audio tokens through modality weaving and strategic global attention.
@@ 4.1.2 Conditional Flow Matching for High-Fidelity Waveforms
The predicted units serve as content for a CFM decoder parameterized by a Diffusion Transformer (DiT), which reconstructs mel-spectrograms; waveforms are then obtained using Vocos (Siuzdak, 2023). Let $\mathbf{x}_{1}$ be the target mel and $\mathbf{x}_{0}\!\sim\!\mathcal{N}(0,\sigma^{2}\mathbf{I})$ a noise prior. Flow matching learns a velocity field $v_{\theta}$ along the linear path $\mathbf{x}_{t}=(1{-}t)\mathbf{x}_{0}+t\mathbf{x}_{1}$.
@@
The condition $\mathbf{c}$ follows a dual-path design: (i) a content stream obtained by embedding unit indices and temporally interpolating them to the mel frame rate; and (ii) a global timbre stream formed by time-averaging BEATs features and broadcasting them across time. During training and validation, we additionally expose a short reference prefix of the ground-truth mel in the conditioning branch (zero-padded to full length) to encourage accurate continuation while keeping the main path free-form.
@@ 4.2.1 Input-Level Modality Weaving
$\mathbf{E}_{[A]}\;\leftarrow\;\mathbf{W}\,\mathrm{Align}\!\big(\Phi_{\text{{BEATs}{}}}(x)\big),\qquad\mathbf{W}\!\in\!\mathbb{R}^{D\times H},\ \mathbf{E}_{[A]}\!\in\!\mathbb{R}^{T\times H}.$ (6)
@@ 4.2.2 Strategic Global Attention
There is mutual attention between tokens in $\mathcal{G}$ and the entire sequence; all others remain local. This pattern preserves linear-time memory while creating cross-modal hubs that support long-range evidence flow: textual symptoms (e.g., “nocturnal dry cough”) can directly query transient acoustic events even when far apart.
@@ 6 Conclusion
We present Resp-Agent, a centrally orchestrated, closed-loop multi-agent framework that unifies controllable, high-fidelity respiratory sound synthesis with multimodal disease diagnosis. Our framework is underpinned by Resp-229k, a large-scale cross-domain benchmark with a strict evaluation protocol. A curriculum-aware Thinker-A2CA planner decomposes diagnostic goals and allocates them between a controllable multimodal Generator and a modality-weaving Diagnoser, turning previously isolated modules into an analyze–synthesize loop. We envision clinician-in-the-loop deployments in which Resp-Agent synthesizes edge-case exemplars and provides audio-informed decision support at the point of care, advancing trustworthy medical-audio AI.
@@
The Resp-229k benchmark is curated from multiple public respiratory-sound corpora, including ICBHI, SPRSound, UK COVID-19, COUGHVID, KAUH, and HF_Lung_V1. Each recording retains explicit provenance, including the source dataset identifier, device metadata (e.g., an electronic stethoscope vs. a microphone), and sampling rate. To ensure rigorous evaluation, we enforce a strict cross-institution and cross-device protocol: the training and validation sets are derived exclusively from ICBHI, SPRSound, and UK COVID-19, whereas the test set consists solely of unseen recordings from KAUH and COUGHVID. All source data were de-identified by their original custodians prior to public release.
@@
We strictly adhere to the original licenses and terms of use for all constituent datasets. Specifically, our usage complies with the Open Government Licence v3.0 (UK COVID-19), Creative Commons Attribution 4.0 International (CC BY 4.0) (COUGHVID, HF_Lung_V1, KAUH, SPRSound), and CC0 Public Domain Dedication (ICBHI). Our derived code, models, and synthetic examples will be released under open-source terms that are compatible with these licenses to facilitate reproducible research while preserving data privacy.
@@ Appendix C Experimental Setup
Split and task. Official ICBHI 60–40% train–test split for four-class RSC (Normal, Crackle, Wheeze, Both); metrics are Specificity (Sp), Sensitivity (Se), and the ICBHI Score $\tfrac{1}{2}(\mathrm{Sp}+\mathrm{Se})$.
Fairness. On ICBHI we use plain cross-entropy with no class reweighting or resampling to match prior practice; reported results follow the official metric definitions.
Optimization. AdamW with a OneCycleLR schedule peaking at $1\times 10^{-5}$; gradient checkpointing is enabled for memory efficiency. Unless otherwise stated, all remaining hyperparameters (token budgets, feature checkpoints, anchor stride) are specified in our released configs and scripts.
Pretraining (summary). We pretrain on HF_Lung_V1 and SPRSound with Focal Loss to emphasize minority pathologies while using an LLM to render heterogeneous metadata into standardized clinical summaries paired with audio. During ICBHI fine-tuning, we remove focal/balancing heuristics for strict apples-to-apples comparisons.

### v3

# arXiv:2602.15909v3[eess.AS] 27 Feb 2026

@@ Abstract
Deep learning-based respiratory auscultation is currently hindered by two fundamental challenges: (i) inherent information loss, as converting signals into spectrograms discards transient acoustic events and clinical context; (ii) limited data availability, exacerbated by severe class imbalance. To bridge these gaps, we present Resp-Agent, an autonomous multimodal system orchestrated by a novel Active Adversarial Curriculum Agent (Thinker-A2CA). Unlike static pipelines, Thinker-A2CA serves as a central controller that actively identifies diagnostic weaknesses and schedules targeted synthesis in a closed loop. To address the representation gap, we introduce a modality-weaving Diagnoser that weaves clinical text with audio tokens via strategic global attention and sparse audio anchors, capturing both long-range clinical context and millisecond-level transients. To address the data gap, we design a flow matching Generator that adapts a text-only Large Language Model (LLM) via modality injection, decoupling pathological content from acoustic style to synthesize hard-to-diagnose samples. As a foundation for this work, we introduce Resp-229k, a benchmark corpus of 229k recordings paired with LLM-distilled clinical narratives. Extensive experiments demonstrate that Resp-Agent consistently outperforms prior approaches across diverse evaluation settings, improving diagnostic robustness under data scarcity and long-tailed class imbalance. Our code and data are available at https://github.com/zpforlove/Resp-Agent.
@@ 1 Introduction
The first challenge is a unimodal representational bottleneck. Audio models often convert signals into mel-spectrograms for image-style CNNs (Bae et al., 2023; He et al., 2024), which discards phase and blurs fine temporal structure, obscuring transient events such as crackles (Paliwal et al., 2011). Conversely, text-only models capture electronic health record (EHR) context but lack objective acoustic evidence, limiting discrimination between conditions with similar narratives but distinct auscultatory patterns. Notably, large-scale health acoustic pretraining (Baur et al., 2024) remains unimodal; without deep multimodal fusion, performance and reliability saturate.
@@ 2 Related Work
Respiratory Sound Classification (RSC) has largely relied on audio-only pipelines trained on small, single-source datasets, which limits out-of-domain generalization. Standard approaches utilize pretrained backbones like PANNs (Kong et al., 2020) and AST (Gong et al., 2021) to mitigate data scarcity. The OPERA benchmark has begun to close the data gap via domain-specific pretraining (Zhang et al., 2024a); however, most systems remain constrained by single-modality supervision and in-distribution evaluation. Our work departs from this paradigm in two key ways. (i) Multimodal fusion. We weave EHR-style textual tokens and acoustic tokens within a long-context Transformer. Unlike RespLLM (Zhang et al., 2024b), which feeds concatenated modality tokens through dense full attention, we introduce Strategic Global Attention with sparse audio anchors, drawing on ideas from efficient Transformers (Beltagy et al., 2020; Zaheer et al., 2020) to route clinical context to transient acoustic events at sub-quadratic cost. We evaluate on source-disjoint splits to stress-test generalization under realistic distribution shifts (Koh et al., 2021) and label imbalance (Johnson and Khoshgoftaar, 2019). (ii) Targeted augmentation. Recent generative models enable high-fidelity audio synthesis (Borsos et al., 2023; Lipman et al., 2023; Liu et al., 2023; Peebles and Xie, 2023), but existing augmentation strategies are typically untargeted, relying on generic perturbations such as SpecAugment (Park et al., 2019) or unconditional generation (Kim et al., 2023). We introduce Resp-Agent, a closed-loop system in which an LLM-based Thinker-A2CA diagnoses model failures and requests condition-controlled synthesis from a flow-matched generator, turning augmentation into a precise instrument for adversarial edge-case creation and distribution balancing.
@@ 3 Resp-229k: A Large-Scale, Multi-Source, Cross-Domain Benchmark
We introduce Resp-229k to address the scarcity of multimodal supervision and the lack of robust cross-domain evaluation in respiratory sound analysis. Unlike existing datasets, Resp-229k provides paired audio with standardized clinical summaries, converting diverse metadata into a format suitable for multimodal modeling. We also establish a strict out-of-domain evaluation protocol to explicitly test model generalization. We aggregate 238,074 raw clips from five public databases; after discarding corrupted or too-short recordings, 229,101 quality-controlled samples remain, spanning 16 classes (15 conditions and 1 control), and are used in all main-paper experiments.
@@ 4 Resp-Agent: An LLM-Orchestrated Loop for Unified Diagnosis and Controllable Synthesis
The overall architecture of Resp-Agent is depicted in Figure 1. Given the paired text–audio supervision and the cross-domain split established by Resp-229k, Resp-Agent is designed as a centrally planned, compute-aware multi-agent system that integrates standalone audio and NLP modules into a closed loop. A compute-efficient planner, Thinker-A2CA (DeepSeek-V3.2-Exp; (Liu et al., 2025)), performs semantic intent parsing and plan–execute tool routing using structured arguments (audio paths, EHR tables, and target classes), enforcing deterministic I/O schemas and emitting concise, instrumented logs. Beyond dispatch, the controller reuses model rationales, error profiles, and calibrated confidence to bias subsequent actions (e.g., targeted synthesis for failure modes), thereby coupling data generation and diagnosis under tight accelerator budgets without compromising coverage or reproducibility. The Thinker coordinates two task-specific agents detailed below: a Generator (Section 4.1) that synthesizes controllable respiratory audio via modality injection and conditional flow matching, and a Diagnoser (Section 4.2) that fuses clinical narratives with audio tokens through modality weaving and strategic global attention.
@@ 4.1.2 Conditional Flow Matching for High-Fidelity Waveforms
The predicted units serve as content for a CFM decoder parameterized by a Diffusion Transformer (DiT), which reconstructs mel-spectrograms; waveforms are then obtained using Vocos (Siuzdak, 2024). Let $\mathbf{x}_{1}$ be the target mel and $\mathbf{x}_{0}\!\sim\!\mathcal{N}(0,\sigma^{2}\mathbf{I})$ a noise prior. Flow matching learns a velocity field $v_{\theta}$ along the linear path $\mathbf{x}_{t}=(1{-}t)\mathbf{x}_{0}+t\mathbf{x}_{1}$.
@@
The condition $\mathbf{c}$ follows a dual-path design: (i) a content stream obtained by embedding unit indices and temporally interpolating them to the mel frame rate; and (ii) a global timbre stream formed by time-averaging BEATs features and broadcasting them across time. During training, we additionally expose a short reference prefix of the ground-truth mel in the conditioning branch (zero-padded to full length) to encourage accurate continuation while keeping the main path free-form.
@@ 4.2.1 Input-Level Modality Weaving
$\mathbf{E}_{[A]}\;\leftarrow\;\mathrm{Align}\!\big(\Phi_{\text{{BEATs}{}}}(x)\big)\,\mathbf{W},\qquad\mathbf{W}\!\in\!\mathbb{R}^{D\times H},\ \mathbf{E}_{[A]}\!\in\!\mathbb{R}^{T\times H}.$ (6)
@@ 4.2.2 Strategic Global Attention
There is mutual attention between tokens in $\mathcal{G}$ and the entire sequence; all others remain local. This pattern preserves linear-time memory while creating cross-modal hubs that support long-range evidence flow. Textual symptoms (e.g., “nocturnal dry cough”) can directly query transient acoustic events even when far apart.
@@ 6 Conclusion
We present Resp-Agent, a centrally orchestrated, closed-loop multi-agent framework that unifies controllable, high-fidelity respiratory sound synthesis with multimodal disease diagnosis. Our framework is underpinned by Resp-229k, a large-scale cross-domain benchmark with a strict evaluation protocol. A curriculum-aware Thinker-A2CA planner decomposes diagnostic goals and allocates them between a controllable multimodal Generator and a modality-weaving Diagnoser, turning previously isolated modules into a closed analyze–synthesize loop. From a methodological standpoint, Resp-Agent instantiates a general principle for data-scarce medical domains: tightly coupling where the model fails with what the generator synthesizes transforms augmentation from a passive remedy into an active, self-correcting training curriculum. We envision clinician-in-the-loop deployments where Resp-Agent synthesizes edge-case exemplars and delivers audio-informed decision support at the point of care, contributing toward trustworthy and equitable medical-audio AI.
@@
The Resp-229k benchmark is curated from multiple public respiratory-sound corpora, including ICBHI, SPRSound, UK COVID-19, COUGHVID, and KAUH. We additionally use HF_Lung_V1 as an external, publicly available pretraining resource for model initialization; it is not included in Resp-229k. Each recording retains explicit provenance, including the source dataset identifier, device metadata (e.g., an electronic stethoscope vs. a microphone), and sampling rate. To ensure rigorous evaluation, we enforce a strict cross-institution and cross-device protocol: the training and validation sets are derived exclusively from ICBHI, SPRSound, and UK COVID-19, whereas the test set consists solely of unseen recordings from KAUH and COUGHVID. All source data were de-identified by their original custodians prior to public release.
@@
We strictly adhere to the original licenses and terms of use for all constituent datasets. Specifically, our usage complies with the Open Government Licence v3.0 (UK COVID-19), Creative Commons Attribution 4.0 International (CC BY 4.0) (COUGHVID, HF_Lung_V1, KAUH, SPRSound), and CC0 Public Domain Dedication (ICBHI) [^1] The ICBHI 2017 Respiratory Sound Database is described as “freely available for research” on its official website (https://bhichallenge.med.auth.gr/ICBHI_2017_Challenge). The CC0 designation follows the license annotation adopted by prior benchmarking work (Zhang et al., 2024a).. Our derived code, models, and synthetic examples have been released under open-source terms that are compatible with these licenses to facilitate reproducible research while preserving data privacy.
@@ Appendix C Experimental Setup
We use the official ICBHI 60–40% train–test split for four-class RSC (Normal, Crackle, Wheeze, Both); metrics are Specificity (Sp), Sensitivity (Se), and the ICBHI Score $\tfrac{1}{2}(\mathrm{Sp}+\mathrm{Se})$. For fairness, on ICBHI we use plain cross-entropy with no class reweighting or resampling to match prior practice, and reported results follow the official metric definitions. Optimization uses AdamW with a OneCycleLR schedule peaking at $1\times 10^{-5}$, and gradient checkpointing is enabled for memory efficiency; unless otherwise stated, all remaining hyperparameters (token budgets, feature checkpoints, anchor stride) are specified in our released configs and scripts. We pretrain on HF_Lung_V1 and SPRSound with Focal Loss to emphasize minority pathologies while using an LLM to render heterogeneous metadata into standardized clinical summaries paired with audio, and during ICBHI fine-tuning we remove focal/balancing heuristics for strict apples-to-apples comparisons.